# Topological signatures of jamming in granular force networks via persistent homology

### v1

@@ 2 Experimental Setup
We construct a birefringent 2D granular ensemble by cutting disk-shaped particles from a photoelastic sheet (Vishay PSM-4). The ensemble consists of a bidisperse system made up of 832 particles (See figure 1(a)), with diameters $D_{1}$ = 15.4 mm and $D_{2}$ = 11 mm, in a $\approx$1:1 concentration ratio in terms of particle number. The ensemble is confined by rails within an area $\approx$ $0.4$ m ×$\times 0.4$ m, where two movable rails are operated by a linear actuator (12v, 2 Amps) to apply biaxial compression at a speed of 0.38 mm/s. The particles rest on a porous sheet that provides air cushioning, minimizing base friction between the particles and the sheet, which ensures that interactions are solely particle-particle. To implement photoelasticity, we construct a brightfield polariscope to leverage the birefringent property of the particles to extract vector contact force information. The camera used for imaging has a resolution of 134.5 microns and captures images at 1 frame per second (fps). This vector contact force data is used to analyze the topological features of the granular ensemble. We make use of PeGS (Photoelastic Grain Solver - an open source software) Kollmer2024jekollmer , which follows a systematic approach to convert the experimental images into force networks as summarized below:
@@ 3 Methodology
  • $\beta_{0}$ zeroth Betti number, captures the number of connected components, indicating groups of particles in force-carrying clusters.
  • $\beta_{1}$ first Betti number, captures the number of loops or holes, highlighting the presence of circuits in the force network.
  • $\gamma_{0}$ the number of homology generators in the persistence diagram of dimension-0 (connected components).
  • $\gamma_{1}$ the number of homology generators in the persistence diagram of dimension-1 (loops).
@@ 4 Results and Discussions
Figure 3 shows the evolution of $\beta_{1}$/#$\#edges$ with increasing packing fraction. We observe that beyond the transition point ($\phi\approx 0.7755$), the number of loops grows more rapidly than the total number of force-bearing contacts (edges). Rather than forming isolated branches, contacts begin to reinforce local rigidity by building nested or overlapping cycles, a hallmark of robust jammed packing, thereby increasing the count of independent cycles in the network. From a topological standpoint, this behavior indicates a denser and more interconnected force network, where loop proliferation is a dominant mechanism of structural evolution. Notice that this transition starts well before $\phi\approx 0.7755$ and around $\phi\approx 0.774$.
@@
Figure 4: Variation of $\gamma_{0}$ with packing fraction $\phi$. After $\phi=0.7773$ (dotted magenta line), the $\gamma_{0}$ value stays saturated within a band of 160±$\pm{10}$.

### v2

@@ 2 Experimental Setup
We construct a birefringent 2D granular ensemble by cutting disk-shaped particles from a photoelastic sheet (Vishay PSM-4). The ensemble consists of a bidisperse system made up of 832 particles (See figure 1(a)), with diameters $D_{1}$ = 15.4 mm and $D_{2}$ = 11 mm, in a $\approx$1:1 concentration ratio in terms of particle number. The ensemble is confined by rails within an area $\approx$ $0.4$ m $\times 0.4$ m, where two movable rails are operated by a linear actuator (12v, 2 Amps) to apply biaxial compression at a speed of 0.38 mm/s. The particles rest on a porous sheet that provides air cushioning, minimizing base friction between the particles and the sheet, which ensures that interactions are solely particle-particle. To implement photoelasticity, we construct a brightfield polariscope to leverage the birefringent property of the particles to extract vector contact force information. The camera used for imaging has a resolution of 134.5 microns and captures images at 1 frame per second (fps). This vector contact force data is used to analyze the topological features of the granular ensemble. We make use of PeGS (Photoelastic Grain Solver - an open source software) Kollmer2024jekollmer , which follows a systematic approach to convert the experimental images into force networks as summarized below:
@@ 3 Methodology
  • $\beta_{0}$ - zeroth Betti number, captures the number of connected components, indicating groups of particles in force-carrying clusters.
  • $\beta_{1}$ - first Betti number, captures the number of loops or holes, highlighting the presence of circuits in the force network.
  • $\gamma_{0}$ - the number of homology generators in the persistence diagram of dimension-0 (connected components).
  • $\gamma_{1}$ - the number of homology generators in the persistence diagram of dimension-1 (loops).
@@ 4 Results and Discussions
Figure 3 shows the evolution of $\beta_{1}$/#edges\#edges with increasing packing fraction. We observe that beyond the transition point ($\phi\approx 0.7755$), the number of loops grows more rapidly than the total number of force-bearing contacts (edges). Rather than forming isolated branches, contacts begin to reinforce local rigidity by building nested or overlapping cycles, a hallmark of robust jammed packing, thereby increasing the count of independent cycles in the network. From a topological standpoint, this behavior indicates a denser and more interconnected force network, where loop proliferation is a dominant mechanism of structural evolution. Notice that this transition starts well before $\phi\approx 0.7755$ and around $\phi\approx 0.774$.
@@
[…]